975 mg of Rocephin (ceftriaxone) must be prepared. The Rocephin is available in a vial of powered drug containing 2 grams. Directions accompanying the drug state: add 19.2 mL of diluent to yield 100mg in 1 mL. How many mL should you withdraw from the vial after reconstituting the drug as directed?
Answer #1

As per direction add 19.2 ml of diluent to yield 100 mg/ml

100 mg in 1 ml

1 mg in =1/100 ml

2000 mg = (1/100) X2000=20 ml

for 975 mg required ml to be withdrawn= (1/100) X975= 9.75 ml

Hence 9.75 ml should be withdrawn from the vial after reconstituting the drug as directed.